OpenStudy (anonymous):

Describe an inverse operation that will undo :Adding - 11 PLEASE HELP ME!

jimthompson5910 (jim_thompson5910):

If you add -11 to a number, you undo it by adding 11 notice how -11 + 11 = 0

OpenStudy (anonymous):

Yess

jimthompson5910 (jim_thompson5910):

So say for example you start with the number 100 Add -11: 100 + (-11) = 89 Then add 11 back: 89 + 11 = 100 you're back to 100. No change has occurred.

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

well said